OpenStudy (anonymous):

Solve: 4/5 y + 2 =6 (I'll try to solve same as you)

OpenStudy (turingtest):

\[\frac{4y}{5}+2=6\]or\[\frac{4}{5y}+2=6\]?

OpenStudy (anonymous):

4 over 5 y plus two equals six

OpenStudy (turingtest):

what di you mean 'I'll solve same as you'? I have the answer, what did you get?

OpenStudy (turingtest):

\[\frac{4}{5y}+2=6\to 4+10y=30y\to20y=4\to y=\frac{4}{20}=\frac{1}{5}\]what of it?

OpenStudy (anonymous):